The wages of 8 men & 6 boys amount to Rs 33. If 4 men earn Rs 4.50 more than 5 boys determine the wages of each man & boy.
Answer by mananth(16946)

let 5 boys earn Rs x
each boy earns x/5
.....
4 men earm x+4.50
each man earns (x+4.50)/4
..
8 men and 6 boys earn Rs. 33
8/4 * (x+4.50) + 6x/5 = 33
2(x+4.50)+6x/5 =33
2x+9+6x/5=33
2x+6x/5=33-9
(10x+6x)/5 = 24
16x= 5*24
x= 5*24/16
5 boys earn 7.50
1 boy earns 7.50/5
=1.50
Man earns (7.50+4.50)/4
12/4
=3
CHECK
8*3+6*1.5=33
..
m.ananth@hotmail.ca
Man earns Rs3 & boy earns Rs.1.50